ONE CAN’T HEAR ORIENTABILITY OF SURFACES

Pierre Bérard and David L. Webb

Abstract. We prove that there exist pairs of compact flat surfaces with boundary which are isospectral for the Laplacian with Neumann boundary conditions; one surface is orientable, while the other is not. These surfaces are not isospectral for the Laplacian with Dirichlet boundary conditions.

Typeset by AMS-TEX